OpenStudy (anonymous):

Solve 7(x – 2) – 9 = 4x + 5 + 3x.

OpenStudy (anonymous):

7x-14-9=4x+5+3x 7x-4x-3x=5+14+9 0=28 here we cannot calculate x

OpenStudy (uri):

7x-14-9=4x+5+3x 7x-4x-3x=5+14+9 0=28.